OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem?

OpenStudy (anonymous):

here are the choices A- b + g = 112 b = 2g + 16 B- b + g = 112 g = 2b + 16 C- b − g = 112 b = 2g − 16 D- None of the systems above would solve this problem.

OpenStudy (anonymous):

B + 16 = 2G B + G = 112 Ans. is C

OpenStudy (anonymous):

It can't be C I think, because it says -16, when in the problem it says 16 more than

OpenStudy (anonymous):

So it would be A or D

OpenStudy (anonymous):

@mrtdmr26

OpenStudy (anonymous):

Really I think the answer is D. The equations are: b+16=2g and b+g=112. The system has not integer answer

OpenStudy (anonymous):

Yea that's what I think too

OpenStudy (anonymous):

B + 16 = 2G B = 2G - 16 Clear or not :)

OpenStudy (anonymous):

I think the answer is A personally, but I do see where you are coming from.

OpenStudy (anonymous):

If i were you i chose C Because that is an equation... :) You can move elements in it...

OpenStudy (anonymous):

But b - g = 112 is subtracting the boys from the girls. The problem states that the boys AND girls combined equals 112. And the 2nd part of the equation says b = g2 - 16, but it says 16 MORE than twice as much than the girls. So that is my thinking.

OpenStudy (anonymous):

I say A.

OpenStudy (anonymous):

My english is not too good but I´ll try to explain. Answer A: the 2nd eq is wrong. the problem says 16 + (more) than 2g and not 16+(plus) 2g. Ans B: Is a similar error. Ans C:All boys and girls are 112. Hope it helps

OpenStudy (anonymous):

Ohh, sorry, i cant seen minus between B and G :(